A market research firm conducts telephone surveys with a 40% historical response rate. What is the probability that in a new sample of 400 telephone numbers, at least 150 individuals will cooperate and respond to the questions? In other words, what is the probability that the sample proportion will be at least 150/400 = .375?
Answer by stanbon(75887)

A market research firm conducts telephone surveys with a 40% historical response rate. What is the probability that in a new sample of 400 telephone numbers, at least 150 individuals will cooperate and respond to the questions? In other words, what is the probability that the sample proportion will be at least 150/400 = .375?
---
z(0.375) = (0.375-0.4)/sqrt[0.4*0.6/sqrt(400)] = -0.2282
-------------------
P(p-hat <= 0.375) = P(z <= -0.2282 = 0.4097
==============================================
Cheers,
Stan H.
==============